You open a book to two facing pages. The product of the two pages is 600. What are the page numbers?
Found 2 solutions by bmckenzie40, checkley77:
Answer by bmckenzie40(10)

the pages are 24 and 25
24*25=600

Answer by checkley77(12844)

X(X+1)=600
X^2+X-600=0
(X+25)(X-24)=0
X-24=0
X=24 ANS. FOR THE LEFT HAND PAGE.
24+1=25 ANS. FOR THE RIGHT HAND PAGE.
PROOF:
24*25=600
600=600